OpenStudy (anonymous):

LeBron goes to store with six dollars and buys four packs of gum. James goes to the store with five dollars and buys two packs of gum. If each pack of gum costs $p$ dollars, what is the expression that represents how much money LeBron and James have together after buying gum?

OpenStudy (texaschic101):

6 - 4p + 5 - 2p = total money left (T) 11 - 6p = T -6p = T - 11 p = (T - 11) / -6 or p = -1/6T + 11/6

OpenStudy (anonymous):

so the answer is ?

OpenStudy (texaschic101):

either one on the bottom....they can be read the same. I am not sure exactly what answer your teacher wants, but they will both work